# On rate-optimal partitioning classification from observable and from privatised data

### v2

@@ Abstract
In this paper we revisit the classical method of partitioning classification and study its convergence rate under relaxed conditions, both for observable (non-privatised) and for privatised data. Let the feature vector $X$ take values in ${\mathbb{R}}^{d}$ and denote its label by $Y$. Previous results on the partitioning classifier worked with the strong density assumption, which is restrictive, as we demonstrate through simple examples. We assume that the distribution of $X$ is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated to a $d_{a}$ dimensional subspace. Here, we study the problem under much milder assumptions: in addition to the standard Lipschitz and margin conditions, a novel characteristic of the absolutely continuous component is introduced, by which the exact convergence rate of the classification error probability is calculated, both for the binary and for the multi-label cases. Interestingly, this rate of convergence depends only on the intrinsic dimension $d_{a}$.
The privacy constraints mean that the data $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ cannot be directly observed, and the classifiers are functions of the randomised outcome of a suitable local differential privacy mechanism. The statistician is free to choose the form of this privacy mechanism, and here we add Laplace distributed noises to the discontinuations of all possible locations of the feature vector $X_{i}$ and to its label $Y_{i}$. Again, tight upper bounds on the rate of convergence of the classification error probability are derived, without the strong density assumption, such that this rate depends on $2\,d_{a}$.
62G08,
@@ 1 Introduction
Classification is one of the fundamental problems of statistics and machine learning, see Devroye, Györfi and Lugosi [4] and Vapnik [16]. It has countless applications from health care and industry to commerce and finance. A significant portion of these applications include sensitive data, for example, about the health or financial situations of the clients involved, which should be anonymised before it can be processed, see Berrett and Butucea [2]. Nevertheless, anonymisation is more involved than just removing the names and the addresses of the clients, as the data can contain several other types of sensitive information. The concept of differential privacy in Dwork et al. [6] provides a rigorous framework to measure the amount of information privacy. Localised privacy dates back to Warner [17], and local differential privacy (LDP) was formally defined by Duchi, Jordan, and Wainwright [5]. On the one hand, LDP helps getting rid of a trusted third party; on the other hand, it generates a framework for managing nonparamteric regression and classification, cf. Berrett and Butucea [2] and Berrett, Györfi and Walk [3]. The LDP mechanism allows processing data even in cases, when the original, raw data should only be seen by its legitimate data holder.
Let $X$ denote a feature vector, which takes values in ${\mathbb{R}}^{d}$, and let $Y$ be its discrete valued label. The goal of classification is to infer label $Y$ from feature vector $X$, given a finite sample of input-output pairs $(X_{i},Y_{i}),\,i=1,\dots,n$. If $Y$ can only take two possible values, e.g., ±$\pm 1$, we talk about binary classification, otherwise we have a multi-label classification problem.
In this paper, we revisit the well-studied partitioning classifier, see Kohler and Krzyżak [10], as it is one of the archetypal classification methods. Our aim is to analyse the convergence rate of its classification error probability under mild assumptions. We investigate both binary and multi-label classification, covering observable and privatised (anonymised) data. In the latter case, we apply Laplace type randomisation which ensures LDP constraints.
@@
Nevertheless, SDA is fairly restrictive as we demonstrate through Examples 1, 2 and 3, for which SDA is not satisfied. In general, the Lipschitz and the margin parameters do not determine the convergence rate of the error probability. While the approximation error (bias) depends only on the Lipschitz and margin parameters, the estimation error is very sensitive to the behaviour of the distribution of the feature vector around the decision boundary. These motivate the introduction of an additional parameter characterizing the relation between the margin and the low density areas. Specifically, we assume that the distribution of $X$ is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated on a $d_{a}$-dimensional subspace. For the absolutely continuous component, we introduce an additional mild condition, called the combined margin and density assumption, for which the true convergence rate of the estimation error can be calculated. The combined margin and density parameter is an appropriate characterisation of the intersection of the low density region and the decision boundary.
Our convergence rate for the binary case, derived without SDA, greatly improves the SDA independent rate of Kohler and Krzyżak [10], see (5), as it is demonstrated on our specific examples with appropriately chosen parameters. Our result ensures the same optimal rate for Example 1 as the SDA dependent rate of Kohler and Krzyżak, that is (6). Hence, the optimal rate can be achieved even without SDA. Moreover, for Example 2 our bound matches the convergence rate one gets by direct calculation of the rate, showing that the bound is tight. Finally, for Example 3, though the deduced SDA independent convergence rate is worse than the SDA dependent rate of (6), but it still improves the previously known SDA independent rate, i.e. (5). Furthermore, interestingly, this convergence rate only depends on the intrinsic dimension $d_{a}$. We also derive tight upper bounds for the convergence rate for the case of privatised data, which rate depends on $2\,d_{a}$, instead of $2\,d$. Both of these bounds are extended to multi-label classification, as well. Our bounds for privatised partitioning classification refute the conjecture that without SDA the convergence rate could be arbitrarily slow. The structure of the paper is as follows. First, in Section 2, we revisit and improve the error probability bounds of partitioning classifiers for observable data, both for binary and multi-labeled setups. Then, in Section 3, we follow an analogous program for privatised partitioning classifiers. Finally, the proofs are presented in Section 4.
@@ 2 Partitioning classification from observable data
In this section, we study the problem for observable (non-privatised) data. First, we give an overview about the core problem of classification, recall the partitioning classification rule, and state our main assumptions under which we quantify the convergence rate both for binary and for multi-label classification.
The standard setup of binary classification is as follows: let the random feature vector $X$ take values in ${\mathbb{R}}^{d}$, and let its label $Y$ be ±$\pm 1$ valued. We denote by $\mu$ the distribution of $X$, that is, $\mu(A)={\mathbb{P}}(X\in A)$ for all measurable sets $A\subseteq{\mathbb{R}}^{d}$. The task of classification is to decide on $Y$ given $X$, i.e., one aims to find a decision function $D$ defined on the range of $X$ such that $D(X)=Y$ with large probability. If $D$ is an arbitrary (measurable) decision function, then its error probability is denoted by
@@
Let ${\cal P}_{h}=\{A_{h,1},A_{h,2},\ldots\}$ be a cubic partition of ${\mathbb{R}}^{d}$ with cubic cells $A_{h,j}$ of volume $h^{d}$ such that $(0,h]^{d}\in{\cal P}_{h}$. The raw data $\mathcal{D}_{n}$ are assumed to be i.i.d. copies of the random vector $(X,Y)$,
@@
$D_{n}(x)=\mbox{sign }\nu_{n}(A_{h,j}),\quad\mbox{if}\;\;x\in A_{h,j}.$
@@
If $X$ is bounded, and the margin and the Lipschitz conditions on $m$ are satisfied, then Kohler and Krzyżak [10] showed for partitioning classification with suitably chosen $h_{n}$ that
@@
In Example 1 the margin condition holds with $\gamma=1$. Thus, in the literature only the poor rate (5) has been proved, which is $n^{-2/5}$. By an easy calculation, for the choice $h_{n}=n^{-1/3}$, we obtain that the true rate is $n^{-2/3}$, which corresponds to the optimal rate (6). This means that there is space for improvement, i.e., we can show the optimal rate such that the SDA is weakened. The interesting feature of Example 2 is that the density tends to infinity at the decision boundary for $-1<\delta<0$, and the rate of convergence is worse than $n^{-2/3}$, see the details after Theorem 2.1. In Example 3 the density tends to zero at the decision boundary for $m$. Motivated by these examples, we introduce an additional parameter characterizing the relation between the margin and the low density areas.
Assume a decomposition
@@
$f_{h}(x)=\frac{\mu_{a}(A_{h,j})}{\lambda_{d_{a}}(A_{h,j})}=\frac{\int_{A_{h,j}}f(z)\lambda_{d_{a}}(dz)}{\lambda_{d_{a}}(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.\vspace{1mm}$ (8)
The function $f_{h}$ is the expectation of the histogram density estimate.
The combined margin and density condition means that there is a $0<h_{0}$ such that for any $0<h<h_{0}$ and for all $0<t,$ we have
$G_{\sqrt{f_{h}}}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}|m(x)|\leq t\}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$ $\leq c_{\sqrt{f}}\cdot t^{\gamma(\sqrt{f})}$ (9)
with constants $0\leq c_{\sqrt{f}}$ and $0\leq\gamma(\sqrt{f})$.
@@
$\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f(x)}|m(x)|\leq t\}}\frac{1}{\sqrt{f(x)}}\mu_{a}(dx)$ $\leq\tilde{c}\cdot t^{\gamma(\sqrt{f})}$
with a $\tilde{c}>0$, then the combined margin and density condition holds.
Under mild regularity conditions, the following theorem contains tight upper bounds on the rate of convergence of the error probability. Interestingly, only the dimension $d_{a}$ of the absolutely continuous component matters.
@@
Assume that $X$ is bounded, $m$ satisfies the Lipschitz condition (2) and the margin condition (3) with $0\leq\gamma$. If, in addition, the combined margin and density condition (9) holds with $0\leq\gamma(\sqrt{f})$, then
$L(D_{n})-L^{*}$ $=O\left(h^{1+\gamma}\right)+O\left((nh^{d_{a}})^{-(1+\min\{1,\gamma,\gamma(\sqrt{f})\})/2}\right).$ (11)
@@
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\left(n^{-\frac{1+\min\{1,\gamma,\gamma(\sqrt{f})\}}{2+d_{a}}}\right).$ (13)
@@
In Example 1 for all $\delta>0$, the margin condition and the combined margin and density condition hold such that $\gamma=1$ and $\gamma\leq\gamma(\sqrt{f})$ because of Lemma 2.2 below. Thus, (13) results in the rate $n^{-2/3}$, which is the same as the optimal rate in (6). In Example 2, one has $\gamma=\delta+1$ and $\gamma(\sqrt{f})\geq 1$, therefore the bound on the rate in (13) is equal to $n^{-(2+\delta)/3}$ for $-1<\delta\leq 0$ and $n^{-2/3}$ for $\delta>0$. For the choice $h_{n}=n^{-1/3}$, by direct calculation of the error probability, one gets the same convergence rate, hence the bound (13) is tight. In Example 3, one has $\gamma=1$ and $\gamma(\sqrt{f})=3/5$ and so $\gamma>\gamma(\sqrt{f})$. Thus, the rate in Theorem 2.1 is $n^{-8/15}$, which is faster than the poor rate in (5), but it is worse than the rate in (6) with $\gamma=1$.
If there is no combined margin and density condition, then it can be interpreted such that the combined margin and density condition holds with $\gamma(\sqrt{f})=0$. Thus, for the choice (12), (13) says that
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\left(n^{-\frac{1}{2+d_{a}}}\right).$
In the bound (11) the first term, called approximation error bound, follows from the Lipschitz condition and from the margin condition, while the second term, called estimation error bound, has been derived from the margin condition and from the combined margin and density condition. Note that the upper bound on the estimation error is managed by CLT approximation with an error term of order $O(1/(nh^{d_{a}}))$. Therefore, the estimation error is upper bounded by $c_{1}\cdot(nh^{d_{a}})^{-(1+\min\{1,\gamma,\gamma(\sqrt{f})\})/2}+c_{0}/(nh^{d_{a}})$, and it is lower bounded by $c^{\prime}_{1}\cdot(nh^{d_{a}})^{-(1+\min\{1,\gamma,\gamma(\sqrt{f})\})/2}-c_{0}/(nh^{d_{a}})$.
@@
and $f_{\epsilon,min}=\inf_{x\in B^{*}_{\epsilon},0<h<h_{0}}f_{h}(x)$. If there is an $10<\epsilon<$ 1 such that $f_{\epsilon,min}>0$, then $\gamma(\sqrt{f})\geq\gamma$, therefore $\gamma=\min\{\gamma,\gamma(\sqrt{f})\}$.
Next, we consider the classification problem for multiple classes. It means that $Y$ takes values in $\{1,\dots,M\}$. Denote by
@@
($k=1,\dots,M$) the a posteriori probabilities. Then, the Bayes decision has the form
@@
Based on ${\cal D}_{n}$, one can estimate the regression functions $P_{k}$ by $P_{n,k}$, and the plug-in classification rule $g_{n}$ derived from $P_{n,k}$ is
@@
For multiple classes, the combined margin and density condition means that there is a $0<h_{0}$ such that for any $0<h<h_{0}$ and for all $0<t,$ we have
$G_{\sqrt{f_{h}}}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}(P_{(1)}(x)-P_{(2)}(x))\leq t\}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$ $\leq c_{\sqrt{f}}\cdot t^{\gamma(\sqrt{f})}$ (15)
with constants $0\leq c_{\sqrt{f}}$ and $0\leq\gamma(\sqrt{f})$.
For multiple classes, the partitioning rule is defined as follows: Put
@@
and
$D_{n}(x)=\mathop{\rm arg\,max}_{k}\nu_{n,k}(A_{h,j}),\quad\mbox{if }x\in A_{h,j}.$
@@
Assume that $X$ is bounded. If $P_{1},\dots,P_{M}$ satisfies the Lipschitz condition (2), the margin condition (14) with $0\leq\gamma$ and the combined margin and density condition (15) with $0\leq\gamma(\sqrt{f})$, then
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h^{1+\gamma}\right)+O\left((nh^{d_{a}})^{-(1+\min\{1,\gamma,\gamma(\sqrt{f})\})/2}\right).$

### v3

@@ Abstract
In this paper we revisit the classical method of partitioning classification and study its convergence rate under relaxed conditions, both for observable (non-privatised) and for privatised data. We consider the problem of classification in a $d$ dimensional Euclidean space. Previous results on the partitioning classifier worked with the strong density assumption, which is restrictive, as we demonstrate through simple examples. Here, we study the problem under much milder assumptions. We presuppose that the distribution of the inputs is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated to a $d_{a}$ dimensional subspace. In addition to the standard Lipschitz and margin conditions, a novel characteristic of the absolutely continuous component is introduced, by which the exact convergence rate of the classification error probability is computed, both for the binary and for the multi-label cases. Interestingly, this rate of convergence depends only on the intrinsic dimension of the inputs, $d_{a}$. The privacy constraints mean that the independent identically distributed data cannot be directly observed, and the classifiers are functions of the randomised outcome of a suitable local differential privacy mechanism. In this paper we add Laplace distributed noises to the discontinuations of all possible locations of the feature vector and to its label. Again, tight upper bounds on the rate of convergence of the classification error probability are derived, without the strong density assumption, such that this rate depends on $2d_{a}$.
62G05,
@@ 1 Introduction
Classification is one of the fundamental problems of statistics and machine learning, see Devroye, Györfi and Lugosi [6] and Vapnik [20]. It has countless applications from health care and industry to commerce and finance. A significant portion of these applications include sensitive data, for example, about the health or financial situations of the clients involved, which should be anonymised before it can be processed. Nevertheless, anonymisation is more involved than just removing the names and the addresses of the clients, as the data can contain several other types of sensitive information. The concept of differential privacy in Dwork et al. [8] provides a rigorous framework to measure the amount of information privacy. Localised privacy dates back to Warner [21], and local differential privacy (LDP) was formally defined by Duchi, Jordan, and Wainwright [7]. On the one hand, LDP helps getting rid of a trusted third party; on the other hand, it generates a framework for managing nonparamteric regression and classification, cf. Berrett and Butucea [3] and Berrett, Györfi and Walk [4]. The LDP mechanism allows processing data even in cases, when the original, raw data should only be seen by its legitimate data holder.
In this paper, we focus on partitioning rules, see Kohler and Krzyżak [14], which are archetypical classification methods. They are key componenents of several machine learning methods, such as decision trees, random forrests, piecewise estimators and hierarchical models. They provide viable solutions to challenges related to scalability, interpretability and explainability, henceforth, they are well-suited for integration into, e.g., federated learning [1] and ensemble-based methods [5]. They are especially useful in low-dimensional settings, provide benchmark models for complex data structures and can be applied for hyperparameter tuning and data cleaning. In general, local averaging estimators are easier to understand compared to complex models and can be used to prove rate of convergence and minimax results in nonparametric settings. The main advantages of the partitioning classifier and its generalizations, like prototype classifiers are their small computational complexity and the sample compression property. Our aim is to analyse the convergence rate of its classification error probability under mild statistical assumptions. We investigate both binary and multi-label classification, covering observable and privatised (anonymised) data. In the latter case, we apply Laplace type randomisation which ensures LDP constraints, cf. [3].
@@
Nevertheless, SDA is fairly restrictive as we demonstrate through Examples 1, 2 and 3, for which SDA is not satisfied. In general, the Lipschitz and the margin parameters do not determine the convergence rate of the error probability. While the approximation error (bias) depends only on the Lipschitz and margin parameters, the estimation error is very sensitive to the behaviour of the distribution of the feature vector around the decision boundary. These motivate the introduction of an additional parameter characterizing the relation between the margin and the low density areas. Specifically, we assume that the distribution of input $X$ is a mixture of an absolutely continuous and a discrete distribution, such that the absolutely continuous component is concentrated on a $d_{a}$-dimensional subspace. For the absolutely continuous component, we introduce an additional mild condition, called the combined margin and density assumption, for which the true convergence rate of the estimation error can be derived. The combined margin and density parameter is an appropriate characterisation of the intersection of the low density region and the decision boundary.
Our convergence rate for the binary case, proved without SDA, greatly improves the SDA independent rate of Kohler and Krzyżak [14], see (5), as it is demonstrated on our specific examples with appropriately chosen parameters. Our result ensures the same optimal rate for Example 1 as the SDA dependent rate of Kohler and Krzyżak, that is (6). Hence, the optimal rate can be achieved even without SDA. Moreover, for Example 2 our bound matches the convergence rate one gets by direct calculation of the rate, showing that the bound is tight. Finally, for Example 3, though the deduced SDA independent convergence rate is worse than the SDA dependent rate of (6), but it still improves the previously known SDA independent rate, that is (5). Furthermore, interestingly, this convergence rate only depends on the intrinsic dimension of the absolutely continuous part, $d_{a}$. We also derive upper bounds for the convergence rate for the case of privatised data with LDP guarantees, which rate depends on $2d_{a}$, instead of $d_{a}$ that was the case for the nonprivate bound. Both of these bounds are extended to multi-label classification, as well. Our bounds for privatised partitioning classification refute the conjecture that without SDA the convergence rate could be arbitrarily slow.
The structure of the paper is as follows. First, in Section 2, we revisit and improve the error probability bounds of partitioning classifiers for observable data, both for binary and multi-labeled setups. Then, in Section 3, we follow an analogous program for privatised partitioning classifiers. The results are summarized and discussed in Section 4 and, finally, the detailed proofs are presented in Section 5.
@@ 2 Partitioning classification from observable data
In this section, we study the problem of classification from observable (non-privatised) data. First, we give an overview about the core problem, recall the partitioning classification rule, and state our main assumptions under which we quantify the convergence rate both for binary and for multi-label classification.
The standard setup of binary classification is as follows: let the random feature vector $X$ take values in ${\mathbb{R}}^{d}$, and let its label $Y$ be $\pm 1$ valued. We denote by $\mu$ the distribution of $X$, that is, $\mu(A)={\mathbb{P}}(X\in A)$ for all measurable sets $A\subseteq{\mathbb{R}}^{d}$. The task of classification is to decide on $Y$ given $X$, i.e., one aims to find a decision function $D$ defined on the range of $X$ such that $D(X)=Y$ with large probability. If $D$ is an arbitrary (measurable) decision function, then its error probability is denoted by
@@
Let ${\cal P}_{h}=\{A_{h,1},A_{h,2},\ldots\}$ be a cubic partition of ${\mathbb{R}}^{d}$ with cubic cells $A_{h,j}$ of volume $h^{d}$ such that $(0,h]^{d}\in{\cal P}_{h}$. The raw data $\mathcal{D}_{n}$ are assumed to be independent identically distributed (i.i.d.) copies of the random vector $(X,Y)$,
@@
$D_{n}(x)=\mbox{sign}(\nu_{n}(A_{h,j})),\quad\mbox{if}\;\;x\in A_{h,j}.$
@@
If $X$ is bounded, and the margin and the Lipschitz conditions on $m$ are satisfied, then Kohler and Krzyżak [14] showed for partitioning classification with suitably chosen $(h_{n})$ that
@@
In Example 1 the margin condition holds with $\gamma=1$. Thus, in the literature only the poor rate (5) has been proved, which is $n^{-2/5}$. By an easy calculation, for the choice $h_{n}=n^{-1/3}$, we obtain that the true rate is $n^{-2/3}$, which corresponds to the optimal rate (6). This means that there is space for improvement, i.e., we can improve the rate if the SDA is weakened. The interesting feature of Example 2 is that the density tends to infinity at the decision boundary for $1<\delta<0$, and the rate of convergence is worse than $n^{-2/3}$, see the details after Theorem 2.1. In Example 3 the density tends to zero at the decision boundary for $m$.
Motivated by these examples, we introduce an additional parameter characterizing the relation between the margin and the low density areas. We assume a decomposition
@@
$f_{h}(x)=\frac{\mu_{a}(A_{h,j})}{\lambda_{d_{a}}(A_{h,j})}=\frac{\int_{A_{h,j}}f(z)\lambda_{d_{a}}(dz)}{\lambda_{d_{a}}(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.\vskip 2.84526pt$ (8)
One can observe that function $f_{h}$ is the expectation of the histogram density estimate.
The combined margin and density condition intuitively controls the measure of a region which is either close to the decision boundary or is such that the density of $X$ is small. Assume that there exists $h_{1}^{*}>0$ such that for any $h\in(0,h_{1}^{*})$ and for all $t>0,$ we have
$G_{h}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}|m(x)|\leq t\}}\frac{f(x)}{\sqrt{f_{h}(x)}}\,\lambda_{d_{a}}\!(dx)$ $\leq c_{1}\cdot t^{\gamma_{1}}$ (9)
with constants $c_{1}>0$ and $0\leq\gamma_{1}=\gamma_{1}(\sqrt{f})$.
@@
$\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f(x)}|m(x)|\leq t\}}\frac{1}{\sqrt{f(x)}}\mu_{a}(dx)$ $\leq\tilde{c}\cdot t^{\gamma_{1}}$
with a $\tilde{c}>0$ and $\gamma_{1}\geq 0$, then the combined margin and density condition holds.
Under mild regularity conditions, the following theorem contains tight upper bounds on the rate of convergence of the error probability. Interestingly, only the dimension $d_{a}$ of the absolutely continuous component matters. The proof is presented in Section 5.
@@
Assume that $X$ is bounded, $m$ satisfies the Lipschitz condition (2) and the margin condition (3) with $0\leq\gamma$. In addition, the combined margin and density condition (9) holds with $\gamma_{1}\geq 0$. Let $(h_{n})$ be a monotonic decreasing null sequence. Then, we have
$L(D_{n})-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left((nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}\right).$ (11)
@@
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\Big{(}n^{-\frac{1+\min\{1,\gamma,\gamma_{1}\}}{2+d_{a}}}\Big{)}.$ (13)
@@
In Example 1 for all $\delta>0$, the margin condition and the combined margin and density condition hold such that $\gamma=1$ and $\gamma\leq\gamma_{1}$ because of Lemma 2.2 below. Thus, (13) results in the rate $n^{-2/3}$, which is the same as the optimal rate in (6). In Example 2, one has $\gamma=\delta+1$ and $\gamma_{1}\geq 1$, therefore the bound on the rate in (13) is equal to $n^{-(2+\delta)/3}$ for $1<\delta\leq 0$ and $n^{-2/3}$ for $\delta>0$. For the choice $h_{n}=n^{-1/3}$, by direct calculation of the error probability, one gets the same convergence rate, hence the bound (13) is tight. In Example 3, one has $\gamma=1$ and $\gamma_{1}=3/5$ and so $\gamma>\gamma_{1}$. Thus, the rate in Theorem 2.1 is $n^{-8/15}$, which is faster than the poor rate in (5), but it is worse than the rate in (6) with $\gamma=1$.
The combined margin and density condition always holds with $\gamma_{1}=0$ in which case for the choice of (12), (13) says that
${\mathbb{E}}\{L(D_{n})\}-L^{*}$ $=O\Big{(}\,n^{-\frac{1}{2+d_{a}}}\,\Big{)}.$
In the bound (11) the first term, called approximation error bound, follows from the Lipschitz condition and from the margin condition, while the second term, called estimation error bound, has been derived from the margin condition and from the combined margin and density condition. Note that the upper bound on the estimation error is managed by CLT approximation with an error term of order $O(1/(nh_{n}^{d_{a}}))$. Therefore, the estimation error is upper bounded by $c_{1}\cdot(nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}+c_{0}/(nh_{n}^{d_{a}})$, and it is lower bounded by $c^{\prime}_{1}\cdot(nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}-c_{0}/(nh_{n}^{d_{a}})$.
@@
and $f_{\epsilon,min}=\inf_{x\in B^{*}_{\epsilon},0<h<h_{0}}f_{h}(x)$. If there is an $\epsilon\in(0,1)$ such that $f_{\epsilon,min}>0$, then $\gamma_{1}\geq\gamma$, therefore $\gamma=\min\{\gamma,\gamma_{1}\}$.
Next, we consider the classification problem for multiple classes. It means that $Y$ takes values in $\{1,\dots,M\}$. Let
@@
denote the a posteriori probabilities for $k=1,\dots,M$. Then, the Bayes decision has the form
@@
Let $P_{n,k}$ be an estimate of $P_{k}$ based on ${\cal D}_{n}$, then the plug-in classification rule $g_{n}$ derived from $P_{n,k}$ is defined by
@@
For multiple classes, the combined margin and density condition means that there is a $h_{1}^{*}>0$ such that for any $h\in(0,h_{1}^{*})$ and for all $0<t,$ we have
$G_{h}(t):=\int_{S_{a}}{\mathbb{I}}_{\{0<\sqrt{f_{h}(x)}(P_{(1)}(x)-P_{(2)}(x))\leq t\}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$ $\leq c_{1}\cdot t^{\gamma_{1}}$ (15)
with constants $0<c_{1}$ and $0\leq\gamma_{1}$.
For multiple classes, the partitioning rule is defined by
@@
and the corresponding plug-in rule as
$D_{n}(x)=\mathop{\rm arg\,max}_{k}\nu_{n,k}(A_{h,j}),\quad\mbox{ for }x\in A_{h,j}.$
Our main result for the multiclass problem is as follows. Its proof is presented in Section 5.
@@
Assume that $X$ is bounded. Additionally, assume that $P_{1},\dots,P_{M}$ satisfy the Lipschitz condition (2), the margin condition (14) with $0\leq\gamma$ and the combined margin and density condition (15) with $\gamma_{1}\geq 0$. Let $(h_{n})$ be a monotonically decreasing sequence with zero limit. Then, we have
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left((nh_{n}^{d_{a}})^{-(1+\min\{1,\gamma,\gamma_{1}\})/2}\right).$